OpenStudy (anonymous):

A circle is intercepted by five central angles each measures m, n, (p + 45), 75, and 111 degrees. Which is m + n + p? Answer 134 degrees 132 degrees 131 degrees 129 degrees

OpenStudy (anonymous):

sum of all central angles has to be 360° 75° + 111° = 186° so m + n + (p + 45) = 360° - 186° m + n + (p + 45) = 174° m + n + p = 174° - 45° m + n + p = 129°